OpenStudy (anonymous):

Which of the following numbers is closest to 0 ? -1/8 1/4 0.2% -|-0.1| π

OpenStudy (anonymous):

-1/8 = -0.125 1/4 = 0.25 0.2% = 0.2/100 = 2/1000 = 0.002 -\(|\)-1\(|\) = -0.1 \(\pi\) = 3.1415...

OpenStudy (anonymous):

so... which one is closest to 0?

OpenStudy (anonymous):

I think its -|-1|

OpenStudy (anonymous):

Or 0.2%

OpenStudy (anonymous):

0.002 is closer to zero, compared to -1, just apply the modulus of teh difference: \(|\)-1-0\(|\) = 1 \(|\)0.002-0\(|\) = 0.002 But 0.002 < 1, so 0.2% is the closest one to zero

OpenStudy (anonymous):

Yeah I agree @M4thM1nd

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

Now the next question is which one is farthest from zero ?

OpenStudy (anonymous):

From those I listed above, which one you think is the larger number?

OpenStudy (anonymous):

π 3.1415

OpenStudy (anonymous):

That's it

OpenStudy (anonymous):

Thanks :) I really appreciated it :)

OpenStudy (anonymous):

No problem ^^